OpenStudy (anonymous):

Which expressions are equal to the expression –16.8 – 4.7? Choose all answers that are correct. A. –4.7 – 16.8 B. –4.7 + 16.8 C. –16.8 + (–4.7) D. 16.8 – (–4.7)

OpenStudy (anonymous):

need severe help

OpenStudy (anonymous):

A and c

OpenStudy (anonymous):

Definitely A

OpenStudy (anonymous):

thank you all :)

OpenStudy (anonymous):

D is not correct tho

OpenStudy (anonymous):

A,C

OpenStudy (anonymous):

ACD

OpenStudy (anonymous):

D isnt an option because 16-(-4.7)=16.8+4.7=21.5

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

@ir.enem exactly